# Tutorial on amortized optimization

### v4

@@ Abstract.
Optimization is a ubiquitous modeling tool and is often deployed in settings which repeatedly solve similar instances of the same problem. Amortized optimization methods use learning to predict the solutions to problems in these settings, exploiting the shared structure between similar problem instances. These methods have been crucial in variational inference and reinforcement learning and are capable of solving optimization problems many orders of magnitudes times faster than traditional optimization methods that do not use amortization. This tutorial presents an introduction to the amortized optimization foundations behind these advancements and overviews their applications in variational inference, sparse coding, gradient-based meta-learning, control, reinforcement learning, convex optimization, optimal transport, and deep equilibrium networks. The source code for this tutorial is available at https://github.com/facebookresearch/amortized-optimization-tutorial.
@@
Figure 1.1: Illustration of the parametric optimization problem in eq. 1.1. Each context $x$ parameterizes an optimization problem that the objective $f(y;x)$ depends on. The contours show the values of the objectives where darker colors indicate higher values. The objective is then minimized over $y$ and the resulting solution $y^{\star}(x)$ is shown in red. In other words, each vertical slice is an optimization problem and this visualization shows a continuum of optimization problems.
@@
where the non-convex objective $f:{\mathcal{Y}}\times{\mathcal{X}}\rightarrow\mathbb{R}$ takes a context or parameterization $x\in{\mathcal{X}}$ which can be continuous or discrete, and the continuous, unconstrained domain of the problem is $y\in{\mathcal{Y}}=\mathbb{R}^{n}$. Eq. 1.1 implicitly defines a solution $y^{\star}(x)\in{\mathcal{Y}}$. In most of the applications considered later in chapter 3, $y^{\star}(x)$ is unique and smooth, i.e., the solution continuously changes in a connected way as the context change s illustrated in fig. 1.1.
@@
Parametric optimization problems such as eq. 1.1 have been studied for decades [Bank et al., 1982, Fiacco and Ishizuka, 1990, Shapiro, 2003, Klatte and Kummer, 2006, Bonnans and Shapiro, 2013, Still, 2018, Fiacco, 2020] with a focus on sensitivity analysis. The general formulation in eq. 1.1 captures many tasks arising in physics, engineering, mathematics, control, inverse modeling, and machine learning. For example, when controlling a continuous robotic system, ${\mathcal{X}}$ is the space of observations or states, e.g., angular positions and velocities describing the configuration of the system, the domain ${\mathcal{Y}}\vcentcolon={\mathcal{U}}$ is the control space, e.g., torques to apply to each actuated joint, and $f(u;x)\vcentcolon=-Q(u,x)$ is the control cost or the negated Q-value of the state-action tuple $(x,u)$, e.g., to reach a goal location or to maximize the velocity. For every encountered state $x$, the system is controlled by solving an optimization problem in the form of eq. 1.1. While ${\mathcal{Y}}=\mathbb{R}^{n}$ is over a deterministic real-valued space in eq. 1.1, the formulation can also capture stochastic optimization problems as discussed in section 2.3.1. For example, Section 3.1 optimizes over the (real-valued) parameters of a variational distribution and section 3.6 optimizes over the (real-valued) parameters of a stochastic policy for control and reinforcement learning.
@@
Figure 1.2: An amortized optimization method learns a model $\hat{y}_{\theta}$ to predict the minimum of an objective $f(y;x)$ to a parameterized optimization problem, as in eq. 1.1, which depends on a context $x$. For example, in control, the context space ${\mathcal{X}}$ is the state space of the system, e.g. angular positions and velocities describing the configuration of the system, the domain ${\mathcal{Y}}\vcentcolon={\mathcal{U}}$ is the control space, e.g. torques to apply to each actuated joint, the cost (or negated value) of a state-action pair is $f(u;x)\vcentcolon=-Q(x,u)$, and the state distribution is $p(x)$. For an encountered state $x$, many reinforcement learning policies $\pi_{\theta}(x)\vcentcolon=\hat{y}_{\theta}(x)$ amortize the solution to the underlying control problem with true solution $y^{\star}(x)$. This humanoid policy was obtained with the model-based stochastic value gradient in Amos et al. [2021].
@@
Optimization problems such as eq. 1.1 quickly become a computational bottleneck in systems they are a part of. These problems often does not have a closed-form analytic solution and is instead solved with approximate numerical methods which iteratively search for the solution. This computational problem has led to many specialized solvers that leverage domain-specific insights to deliver fast solves. Specialized algorithms are especially prevalent in convex optimization methods for linear programming, quadratic programming, cone programming, and control and use theoretical insights of the problem structure to bring empirical gains of computational improvements and improved convergence [Boyd et al., 2004, Nocedal and Wright, 2006, Bertsekas, 2015, Bubeck et al., 2015, Nesterov et al., 2018].
@@
Overview. This tutorial studies the use of machine learning models to rapidly predict the solutions to the optimization problem in eq. 1.1, which is referred to as amortized optimization or learning to optimize. Amortized optimization methods are capable of significantly improving the computational time of classical algorithms on a focused subset of problems. This is because the model is able to learn about the solution mapping from $x$ to $y^{\star}(x)$ that classical optimization methods usually do not assume access to. My goal in writing this is to explore a unified perspective of modeling approaches of amortized optimization in chapter 2 to help draw connections between the applications in chapter 3, e.g. between amortized variational inference, meta-learning, and policy learning for control and reinforcement learning, sparse coding, convex optimization, optimal transport, and deep equilibrium networks. These topics have historically been studied in isolation without connections between their amortization components. Chapter 4 presents a computational tour through source code for variational inference, policy learning, and a spherical optimization problem and chapter 5 concludes with a discussion of challenges, limitations, open problems, and related work.
@@
How much does amortization help? Amortized optimization has been revolutionary to many fields, especially including variational inference and reinforcement learning. Figure 4.1 shows that the amortization component of a variational autoencoder trained on MNIST is 25000 times faster (0.4ms vs. 8 seconds!) than solving a batch of 1024 optimization problems from scratch to obtain a solution of the same quality. These optimization problems are solved in every training iteration and can become a significant bottleneck if they are inefficiently solved. If the model is being trained for millions of iterations, then the difference between solving the optimization problem in 0.4ms vs. 8 seconds makes the difference between the entire training process finishing in a few hours or a month.
@@
A historic note: amortization in control and statistical inference. Amortized optimization has arisen in many fields as a result to practical optimization problems being non-convex and not having easily computed, or closed-form solutions. Continuous control problems with linear dynamics and quadratic cost are convex and often easily solved with the linear quadratic regulator (LQR) and many non-convex extensions and iterative applications of LQR have been successful over the decades, but becomes increasingly infeasible on non-trivial systems and in reinforcement learning settings where the policy often needs to be rapidly executed. For this reason, the reinforcement learning community almost exclusively amortizes control optimization problems with a learned policy [Sutton and Barto, 2018]. Related to this throughline in control and reinforcement learning, many statistical optimization problems have closed form solutions for known distributions such as Gaussians. For example, the original Kalman filter is defined with Gaussians and the updates take closed-form. The extended Kalman filter generalizes the distributions to non-Gaussians, but the updates are in general no longer available analytically and need to be computationally estimated. Marino et al. [2018a] shows how amortization helps improve this computationally challenging step. Both of these control and statistical settings start with a simple setting with analytic solutions to optimization problems, generalize to more challenging optimization problems that need to be computationally estimated, and then add back some computational tractability with amortized optimization.
@@
Amortization model $\hat{y}_{\theta}$ (section 2.1) Fully-amortized (section 2.1.1): no objective access Semi-amortized (section 2.1.2): accesses objective Amortization loss ${\mathcal{L}}$ (section 2.2) Regression (section 2.2.1): $\operatorname*{\mathbb{E}}_{p(x)}\|\hat{y}_{\theta}(x)-y^{\star}(x)\|_{2}^{2}$ Objective (section 2.2.1): $\operatorname*{\mathbb{E}}_{p(x)}f(\hat{y}_{\theta}(x))$
@@
The machine learning, statistics, and optimization communities are exploring methods of learning to optimize to obtain fast solvers for eq. 1.1. I will refer to these methods as amortized optimization as they amortize the cost of solving the optimization problems across many contexts to approximate the solution mapping $y^{\star}$. Amortized optimization is promising because in many applications, there are significant correlations and structure between the solutions which show up in $y^{\star}$ that a model can learn. This tutorial follows Shu [2017] for defining the core foundation of amortized optimization.
@@
An amortized optimization method to solve eq. 1.1 can be represented by ${\mathcal{A}}\vcentcolon=(f,{\mathcal{Y}},{\mathcal{X}},p(x),\hat{y}_{\theta},{\mathcal{L}})$, where $f:{\mathcal{Y}}\times{\mathcal{X}}\rightarrow\mathbb{R}$ is the unconstrained objective to optimize, ${\mathcal{Y}}$ is the domain, ${\mathcal{X}}$ is the context space, $p(x)$ is the probability distribution over contexts to optimize, $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is the amortization model parameterized by $\theta$ which is learned by optimizing a loss defined on all the components ${\mathcal{L}}(f,{\mathcal{Y}},{\mathcal{X}},p(x),\hat{y}_{\theta})$.
@@
The objective $f$ and domain ${\mathcal{Y}}$ arise from the problem setting along with the context space ${\mathcal{X}}$ and distribution over it $p(x)$, and the remaining definitions of the model $\hat{y}_{\theta}$ and loss ${\mathcal{L}}$ are application-specific design decisions that sections 2.1 and 2.2 opens up. These sections present the modeling and loss foundations for the core problem in definition 1 agnostic of specific downstream applications that will use them. The key choices highlighted in fig. 2.1 are how much information 1) the model $\hat{y}_{\theta}$ has about the objective $f$ (fully- vs. semi-amortized), and 2) the loss has about the true solution $y^{\star}$ (regression- vs. objective-based). Figure 1.2 instantiates these components for amortizing the control of a robotic system. The model $\hat{y}_{\theta}$ solves the solution mapping $y^{\star}$ simultaneously for all contexts. The methods here usually assume the solution mapping $y^{\star}$ to be almost-everywhere smooth and well-behaved. The best modeling approach is an open research topic as there are many tradeoffs, and many specialized insights from the application domain can significantly improve the performance. The generalization capacity along with the model’s convergence guarantees are challenging topics which section 5.2 covers in more detail.
@@
Origins of the term “amortization” for optimization. The word “amortization” generally means to spread out costs and thus “amortized optimization” usually means to spread out computational costs of the optimization process. The term originated in the variational inference community for inference optimization [Kingma and Welling, 2014, Rezende et al., 2014, Stuhlmüller et al., 2013, Gershman and Goodman, 2014, Webb et al., 2018, Ravi and Beatson, 2019, Cremer et al., 2018, Wu et al., 2020], and is used more generally in Xue et al. [2020], Sercu et al. [2021], Xiao et al. [2021]. Marino [2021, p. 28] give further background on the origins and uses of amortization. Concurrent to these developments, other communities have independently amortization methods without referring to them by the same terminology and analysis, such as in reinforcement learning, policy optimization, and sparse coding — chapter 3 connects all of these under definition 1.
@@
Conventions and notation. The context space ${\mathcal{X}}$ represents the sample space of a probability space that the distribution $p(x)$ is defined on, assuming it is Borel if not otherwise specified. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ in standard Euclidean space, $\nabla_{x}f(\bar{x})\in\mathbb{R}^{n}$ denotes the gradient at a point $\bar{x}$ and $\nabla^{2}_{x}f(\bar{x})\in\mathbb{R}^{n\times n}$ denotes the Hessian. For $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $\mathrm{D}_{x}f(\bar{x})\in\mathbb{R}^{m\times n}$ represents the Jacobian at $\bar{x}$ with entries $[\mathrm{D}_{x}f(\bar{x})]_{ij}\vcentcolon=\frac{\partial f_{i}}{\partial x_{j}}(\bar{x})$. I abbreviate the loss to ${\mathcal{L}}(\hat{y}_{\theta})$ when the other components can be inferred from the surrounding text and prefer the term “context” for $x$ instead of “parameterization” to make the distinction between the $x$-parameterized optimization problem and the $\theta$-parameterized model clear. I use “;” as separation in $f(y;x)$ to emphasize the separation between the domain variables $y$ that eq. 1.1 optimizes over from the context ones $x$ that remain fixed. A model’s parameters $\theta$ are usually subscripts as $h_{\theta}(x)$ but I will equivalently write $h(x;\theta)$ sometimes.

### v5

@@ Abstract.
Optimization is a ubiquitous modeling tool and is often deployed in settings which repeatedly solve similar instances of the same problem. Amortized optimization methods use learning to predict the solutions to problems in these settings, exploiting the shared structure between similar problem instances. These methods have been crucial in variational inference and reinforcement learning and are capable of solving optimization problems many orders of magnitude faster than traditional optimization methods that do not use amortization. This tutorial presents an introduction to the amortized optimization foundations behind these advancements and overviews their applications in variational inference, sparse coding, gradient-based meta-learning, control, reinforcement learning, convex optimization, optimal transport, and deep equilibrium networks. The source code for this tutorial is available at https://github.com/facebookresearch/amortized-optimization-tutorial.
@@
Figure 1.1: Illustration of the parametric optimization problem in eq.˜$1.$ Each context $x$ parameterizes an optimization problem that the objective $f(y;x)$ depends on. The contours show the values of the objectives where darker colors indicate higher values. The objective is then minimized over $y$ and the resulting solution $y^{\star}(x)$ is shown in red. In other words, each vertical slice is an optimization problem and this visualization shows a continuum of optimization problems.
@@
where the non-convex objective $f:{\mathcal{Y}}\times{\mathcal{X}}\rightarrow\mathbb{R}$ takes a context or parameterization $x\in{\mathcal{X}}$ which can be continuous or discrete, and the continuous, unconstrained domain of the problem is $y\in{\mathcal{Y}}=\mathbb{R}^{n}$. Eq.˜1.1 implicitly defines a solution $y^{\star}(x)\in{\mathcal{Y}}$. In most of the applications considered later in chapter˜3, $y^{\star}(x)$ is unique and smooth, i.e., the solution continuously changes in a connected way as the context changes, as illustrated in fig.˜1.1.
@@
Parametric optimization problems such as eq.˜1.1 have been studied for decades [Bank et al., 1982, Fiacco and Ishizuka, 1990, Shapiro, 2003, Klatte and Kummer, 2006, Bonnans and Shapiro, 2013, Still, 2018, Fiacco, 2020] with a focus on sensitivity analysis. The general formulation in eq.˜1.1 captures many tasks arising in physics, engineering, mathematics, control, inverse modeling, and machine learning. For example, when controlling a continuous robotic system, $\mathcal{X}$} is the space of observations or states, e.g., angular positions and velocities describing the configuration of the system, the domain ${\mathcal{Y}}\vcentcolon={\mathcal{U}}$ is the control space, e.g., torques to apply to each actuated joint, and $f(u;x)\vcentcolon=-Q(u,x)$ is the control cost or the negated Q-value of the state-action tuple $(x,u)$, e.g., to reach a goal location or to maximize the velocity. For every encountered state $x$, the system is controlled by solving an optimization problem in the form of eq.˜1.1. While ${\mathcal{Y}}=\mathbb{R}^{n}$ is over a deterministic real-valued space in eq.˜1.1, the formulation can also capture stochastic optimization problems as discussed in section˜2.3.1. For example, Section˜3.1 optimizes over the (real-valued) parameters of a variational distribution and section˜3.6 optimizes over the (real-valued) parameters of a stochastic policy for control and reinforcement learning.
@@
Figure 1.2: An amortized optimization method learns a model $\hat{y}_{\theta}$ to predict the minimum of an objective $f(y;x)$ to a parameterized optimization problem, as in eq.˜1.1, which depends on a context $x$. For example, in control, the context space $\mathcal{X}$} is the state space of the system, e.g. angular positions and velocities describing the configuration of the system, the domain ${\mathcal{Y}}\vcentcolon={\mathcal{U}}$ is the control space, e.g. torques to apply to each actuated joint, the cost (or negated value) of a state-action pair is $f(u;x)\vcentcolon=-Q(x,u)$, and the state distribution is $p(x)$. For an encountered state $x$, many reinforcement learning policies $\pi_{\theta}(x)\vcentcolon=\hat{y}_{\theta}(x)$ amortize the solution to the underlying control problem with true solution $y^{\star}(x)$. This humanoid policy was obtained with the model-based stochastic value gradient in Amos et al. [2021].
@@
Optimization problems such as eq.˜1.1 quickly become a computational bottleneck in systems they are a part of. These problems often do not have a closed-form analytic solution and are instead solved with approximate numerical methods which iteratively search for the solution. This computational problem has led to many specialized solvers that leverage domain-specific insights to deliver fast solves. Specialized algorithms are especially prevalent in convex optimization methods for linear programming, quadratic programming, cone programming, and control and use theoretical insights of the problem structure to bring empirical gains of computational improvements and improved convergence [Boyd et al., 2004, Nocedal and Wright, 2006, Bertsekas, 2015, Bubeck et al., 2015, Nesterov et al., 2018].
@@
Overview. This tutorial studies the use of machine learning models to rapidly predict the solutions to the optimization problem in eq.˜1.1, which is referred to as amortized optimization or learning to optimize. Amortized optimization methods are capable of significantly improving the computational time of classical algorithms on a focused subset of problems. This is because the model is able to learn about the solution mapping from $x$ to $y^{\star}(x)$ that classical optimization methods usually do not assume access to. My goal in writing this is to explore a unified perspective of modeling approaches of amortized optimization in chapter˜2 to help draw connections between the applications in chapter˜3, e.g. between amortized variational inference, meta-learning, and policy learning for control and reinforcement learning, sparse coding, convex optimization, optimal transport, and deep equilibrium networks. These topics have historically been studied in isolation without connections between their amortization components. Chapter˜4 presents a computational tour through source code for variational inference, policy learning, and a spherical optimization problem and chapter˜5 concludes with a discussion of challenges, limitations, open problems, and related work.
@@
How much does amortization help? Amortized optimization has been revolutionary to many fields, especially including variational inference and reinforcement learning. Figure˜4.1 shows that the amortization component of a variational autoencoder trained on MNIST is 25000 times faster (0.4ms vs. 8 seconds!) than solving a batch of 1024 optimization problems from scratch to obtain a solution of the same quality. These optimization problems are solved in every training iteration and can become a significant bottleneck if they are inefficiently solved. If the model is being trained for millions of iterations, then the difference between solving the optimization problem in 0.4ms vs. 8 seconds makes the difference between the entire training process finishing in a few hours or a month.
@@
A historic note: amortization in control and statistical inference. Amortized optimization has arisen in many fields as a result to practical optimization problems being non-convex and not having easily computed, or closed-form solutions. Continuous control problems with linear dynamics and quadratic cost are convex and often easily solved with the linear quadratic regulator (LQR) and many non-convex extensions and iterative applications of LQR have been successful over the decades, but becomes increasingly infeasible on non-trivial systems and in reinforcement learning settings where the policy often needs to be rapidly executed. For this reason, the reinforcement learning community almost exclusively amortizes control optimization problems with a learned policy [Sutton and Barto, 2018]. Related to this throughline in control and reinforcement learning, many statistical optimization problems have closed form solutions for known distributions such as Gaussians. For example, the original Kalman filter is defined with Gaussians and the updates take a closed form. The extended Kalman filter generalizes the distributions to non-Gaussians, but the updates are in general no longer available analytically and need to be computationally estimated. Marino et al. [2018a] shows how amortization helps improve this computationally challenging step. Both of these control and statistical settings start with a simple setting with analytic solutions to optimization problems, generalize to more challenging optimization problems that need to be computationally estimated, and then add back some computational tractability with amortized optimization.
@@
Amortization model $\hat{y}_{\theta}$ (section˜2.1) Fully-amortized (section˜2.1.1): no objective access Semi-amortized (section˜2.1.2): accesses objective Amortization loss $\mathcal{L}$} (section˜2.2) Regression (section˜2.2.1): $\operatorname*{\mathbb{E}}_{p(x)}\|\hat{y}_{\theta}(x)-y^{\star}(x)\|_{2}^{2}$ Objective (section˜2.2.1): $\operatorname*{\mathbb{E}}_{p(x)}f(\hat{y}_{\theta}(x))$
@@
The machine learning, statistics, and optimization communities are exploring methods of learning to optimize to obtain fast solvers for eq.˜1.1. I will refer to these methods as amortized optimization as they amortize the cost of solving the optimization problems across many contexts to approximate the solution mapping $y^{\star}$. Amortized optimization is promising because in many applications, there are significant correlations and structure between the solutions which show up in $y^{\star}$ that a model can learn. This tutorial follows Shu [2017] for defining the core foundation of amortized optimization.
@@
An amortized optimization method to solve eq.˜1.1 can be represented by ${\mathcal{A}}\vcentcolon=(f,{\mathcal{Y}},{\mathcal{X}},p(x),\hat{y}_{\theta},{\mathcal{L}})$, where $f:{\mathcal{Y}}\times{\mathcal{X}}\rightarrow\mathbb{R}$ is the unconstrained objective to optimize, $\mathcal{Y}$} is the domain, $\mathcal{X}$} is the context space, $p(x)$ is the probability distribution over contexts to optimize, $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is the amortization model parameterized by $\theta$ which is learned by optimizing a loss defined on all the components ${\mathcal{L}}(f,{\mathcal{Y}},{\mathcal{X}},p(x),\hat{y}_{\theta})$.
@@
The objective $f$ and domain $\mathcal{Y}$} arise from the problem setting along with the context space $\mathcal{X}$} and distribution over it $p(x)$, and the remaining definitions of the model $\hat{y}_{\theta}$ and loss $\mathcal{L}$} are application-specific design decisions that sections˜2.1 and 2.2 opens up. These sections present the modeling and loss foundations for the core problem in definition˜1 agnostic of specific downstream applications that will use them. The key choices highlighted in fig.˜2.1 are how much information 1) the model $\hat{y}_{\theta}$ has about the objective $f$ (fully- vs. semi-amortized), and 2) the loss has about the true solution $y^{\star}$ (regression- vs. objective-based). Figure˜1.2 instantiates these components for amortizing the control of a robotic system. The model $\hat{y}_{\theta}$ solves the solution mapping $y^{\star}$ simultaneously for all contexts. The methods here usually assume the solution mapping $y^{\star}$ to be almost-everywhere smooth and well-behaved. The best modeling approach is an open research topic as there are many tradeoffs, and many specialized insights from the application domain can significantly improve the performance. The generalization capacity along with the model’s convergence guarantees are challenging topics which section˜5.2 covers in more detail.
@@
Origins of the term “amortization” for optimization. The word “amortization” generally means to spread out costs and thus “amortized optimization” usually means to spread out computational costs of the optimization process. The term originated in the variational inference community for inference optimization [Kingma and Welling, 2014, Rezende et al., 2014, Stuhlmüller et al., 2013, Gershman and Goodman, 2014, Webb et al., 2018, Ravi and Beatson, 2019, Cremer et al., 2018, Wu et al., 2020], and is used more generally in Xue et al. [2020], Sercu et al. [2021], Xiao et al. [2021]. Marino [2021, p. 28] give further background on the origins and uses of amortization. Concurrent to these developments, other communities have independently developed amortization methods without referring to them by the same terminology and analysis, such as in reinforcement learning, policy optimization, and sparse coding — chapter˜3 connects all of these under definition˜1.
@@
Conventions and notation. The context space $\mathcal{X}$} represents the sample space of a probability space that the distribution $p(x)$ is defined on, assuming it is Borel if not otherwise specified. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ in standard Euclidean space, $\nabla_{x}f(\bar{x})\in\mathbb{R}^{n}$ denotes the gradient at a point $\bar{x}$ and $\nabla^{2}_{x}f(\bar{x})\in\mathbb{R}^{n\times n}$ denotes the Hessian. For $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $\mathrm{D}_{x}f(\bar{x})\in\mathbb{R}^{m\times n}$ represents the Jacobian at $\bar{x}$ with entries $\mathrm{D}_{x}f(\bar{x})]_{ij}\vcentcolon=\frac{\partial f_{i}}{\partial x_{j}}(\bar{x})$. I abbreviate the loss to ${\mathcal{L}}(\hat{y}_{\theta})$ when the other components can be inferred from the surrounding text and prefer the term “context” for $x$ instead of “parameterization” to make the distinction between the $x$-parameterized optimization problem and the $\theta$-parameterized model clear. I use “;” as separation in $f(y;x)$ to emphasize the separation between the domain variables $y$ that eq.˜1.1 optimizes over from the context ones $x$ that remain fixed. A model’s parameters $\theta$ are usually subscripts as $h_{\theta}(x)$ but I will equivalently write $h(x;\theta)$ sometimes.